# Instance-level quantitative saliency in multiple sclerosis lesion segmentation

### v2

@@ Abstract
In recent years, explainable methods for artificial intelligence (XAI) have tried to reveal and describe models’ decision mechanisms in the case of classification tasks. However, XAI for semantic segmentation and in particular for single instances has been little studied to date. Understanding the process underlying automatic segmentation of single instances is crucial to reveal what information was used to detect and segment a given object of interest (e.g., a lesion in medical imaging or a car in a natural image) among others of the same class. This information can serve as a basis to (i) identify application contexts (e.g., lesion location) where the networks may deliver sub-optimal performance, (ii) optimize their internal architecture (e.g., patch size), and (iii) justify the model’s decisions to the end users.
In this study, we proposed two instance-level explanation maps for semantic segmentation based on SmoothGrad and Grad-CAM++ methods. Then, we investigated their relevance for the detection and segmentation of white matter lesions (WML), a magnetic resonance imaging (MRI) biomarker in multiple sclerosis (MS). 687 patients diagnosed with MS for a total of 4043 FLAIR and MPRAGE MRI scans (baseline, follow-ups and WM lesion masks annotated by three expert clinicians) were collected at the University Hospital of Basel, Switzerland. Data were randomly split into training, validation and test sets (containing respectively 585, 94 and 35 patients) to train a 3D U-Net for MS lesion segmentation. With a Dice score of 0.60 and a normalized Dice score of 0.71, we observed 3050 true positive (TP), 1818 false positive (FP), and 789 false negative (FN) cases. We generated instance-level explanation maps for semantic segmentation, by developing two XAI methods based on SmoothGrad and Grad-CAM++. The extent of information used by the U-Net during inference was investigated with the two new instance-based XAI methods at different levels: 1) the distribution of gradients in saliency maps with respect to both input MRI sequences; 2) a comparison between TP, FP, FN and true negative (TN) volumes; 3) the model’s response and saliency map in the case of synthetic lesions in various image locations (in the background, in the WM, and in the background with part of the perilesional tissue); 4) the amount of perilesional tissue needed by the model to segment a lesion.
Saliency maps (based on SmoothGrad) in FLAIR showed positive values inside a lesion and negative in its neighborhood. FLAIR input images consistently had a higher impact on predictions than MPRAGE in TP, FN, FP and TN volumes. Peak values of saliency maps generated for these four groups of volumes presented distributions that differ significantly from one another, suggesting that the quantitative nature of the proposed saliency could be used to identify errors. Artificially isolated lesions (i.e., surrounded by a black background) were not detected as lesions. Contextual information of $\sim 7mm$ around the lesion border was required for their segmentation.
In conclusion, the two proposed quantitative methods provide quantitative instance-level explanations in semantic segmentation tasks, revealing fundamental insights into the decision process of a deep learning MS lesion segmentation network.
@@
[…]
journal: Medical Image Analysis
1 Introduction
Multiple sclerosis (MS) is an autoimmune neurological disease, which affects people at a relatively young age, presenting a considerable impact on the quality of life [16]. One of the most important biomarkers in MS are white matter (WM) lesions reported on magnetic resonance imaging (MRI) [39]. The standard MRI sequences used for MS diagnosis and follow-up [32] are the fluid attenuated inversion recovery (FLAIR) and T1-weighted (T1-w) contrast, such as the magnetisation-prepared rapid gradient echo (MPRAGE) [11]. WM lesions appear as hyperintense in FLAIR and as hypointense in T1-w images.
@@ 1 Introduction
    The“black box” nature of the models [4]. Since these methods contain many layers and millions of parameters it is hard to interpret and explain which are the drivers for a particular decision, i.e., which rules were used to identify and segment a given lesion of interest.
@@
    Insufficient clinical validation of the models, described in [31].
To address the first issue, research in explainable AI (XAI) may play a decisive role. XAI has the potential to support trustworthy AI via better understanding (e.g., decision rules, biases) and optimization of DL models [14]. However, XAI for semantic segmentation and in particular for detecting and contouring single instances of interest has been little studied to date. Semantic segmentation is a computer vision task, where labels are associated with every pixel of an image. In the case of WM lesions, such as depicted in Fig.1a with four distinct lesion instances, the segmented plaques appear as disconnected volumes in the MRI. In general, objects may appear as connected, and can be segmented into separate entities through instance segmentation. Instance segmentation can extract supplementary information from the image, such as the number of objects of the same class (as person one to five in Fig.1b). Treating separate instances would be crucial to understand the mechanisms underlying automatic detection and segmentation of a given object of interest (e.g., a lesion in medical imaging or a person in a natural image) among other objects of the same class. This also applies to the case of MS lesions, where plaques can be subdivided based on their location in the brain, or the stage of the disease [12]. Specifically, considering separate instances would be important to generate instance-specific explanations not only for diagnosis at single time-points, but also for disease monitoring in follow-ups.
@@
Fig. 1: (a) FLAIR MRI presenting WM lesions segmented as separate entities, and (b) example of instance segmentation retrieved from [35].
1.1 Related work
1.1.1 Classification task
@@
An exhaustive review of XAI models and applications can be found in [25]. For convolutional neural networks (CNN) in a classification scenario, a widely used ad-hoc method is pixel attribution (or saliency maps), in which pixels are colored based on their contribution to the classification. To this end, vanilla gradients [28] is a method based on forward and backward propagation through the network. First, an input image (for instance a 3D image) is fed forward through the network and a class score is computed. Then, the gradient of the score with respect to the input — or a layer — is calculated backwards to form a map, which represents positive and negative contributions of input voxels to the classification of the image. Maps generated with this method are easy to compute and visualize, but also noisy and sensitive to small changes in the input [37]. To partially address these problems, [30] introduced a method called SmoothGrad (SG). A more stable output is obtained by feeding multiple noisy versions of the input image to the network and averaging the obtained saliency maps. Some recent applications can be found in [10] and [1].
@@
Another widely used XAI method for classification, is Grad-CAM [26]. The gradients of a class score — with respect to activation maps of a given layer — are spatially aggregated through global average pooling. This way, a weight is computed for each activation map, representing its relevance. These weights are then employed to calculate the visualization heatmap as a linear combination of the activation maps, followed by a ReLU to focus on features positively influencing the class score. However, this method presented a low accuracy when the image contained multiple instances of the same class. To overcome this, [5] described Grad-CAM++, where the weights are obtained through a weighted average of the gradients.
@@
SG and Grad-CAM++ can be seen as complementary methods. The first provides local-level information and works well in identifying the impact of multiple input channels (or modalities), since the gradients flow all the way back to the inputs. The second can generate more stable heatmaps, and probe specific layers of the network. However, both algorithms were originally designed for classification tasks, such as in [24]. This means that, in general, the output $y$ of a multi-class classification network is a vector of $C$ scalar values, where $C$ is the number of classes. In the case of semantic segmentation, $y$ is a set of $C$ tensors (2D or 3D images) containing segmentation scores for each class (Fig. 2).
@@
Fig. 2: Input-output dimensions for a classification (top) and a semantic segmentation (bottom) task.
@@
1.1.2 Segmentation task
A straightforward way to adapt explainable methods to semantic segmentation would be, for a given class $c$, to aggregate all the spatial predictions $y^{c}$ into a single scalar (e.g., a summation) and compute its gradients. However, this approach yields to confusing and hardly interpretable maps, where the relevance of input voxels to the segmentation scores of all output voxels (even those not segmented as part of class $c$) are merged together, as reported in Fig. 3. Indeed, these results do not provide any kind of explanation to the segmentation of a particular instance of the considered class $c$.
Fig. 3: (a) The output of a semantic segmentation network showing several instances of the considered class. SmoothGrad (b) and Grad-CAM (c) applied to all the spatial predictions. (d) How can we explain the segmentation of a particular lesion of interest (e.g., the yellow instance) ?
Recent works, such as [19] and [3], state clearly that saliency maps were not initially developed for segmentation tasks and, therefore, no explicit methods are currently available to do so. In [3], this statement was used to warn the readers about potential problems related to the misuse of XAI in clinical practice, and proposed pixel-level metrics to evaluate saliency maps. [29] applied saliency maps to classification networks, while adopting uncertainty maps instead to explain skin lesion segmentation.
First steps towards the use of XAI in segmentation were taken in the work of [38]. They used guided backpropagation [28] to generate saliency maps for the explanation of colorectal polyp segmentation. Their saliency maps were obtained by aggregating all the positive spatial predictions. Similarly, [36] generated heatmaps by using the first version of Grad-CAM [26], and aggregating only output voxels segmented as part of a target class $c$. The authors employed a U-Net and the dataset Cityscapes [7] to perform semantic segmentation.
@@ 1.1.2 Segmentation task
2 Material and methods
2.1 Dataset and model
@@
687 patients diagnosed with MS for a total of 4023 FLAIR and MPRAGE MRI scans (age=45.2±12.2, 433 females, SIEMENS Avanto/Espree/Symphony 1.5T and Prisma/Skyra/Verio/MAGNETOM Vida 3T, $1mm$ isotropic, Expanded Disability Status Scale median of 2.5 [0-9]) were collected at the University Hospital of Basel, Switzerland. Three expert clinicians annotated WM lesion masks from baseline and follow-ups. Data were randomly split into training, validation and test sets (containing 560, 90 and 37 patients with 3369, 553 and 101 scans, respectively; training/validation set’s and test set’s mean lesions number of 52.9±36.4 and 42.3±21.4 per patient) to train and evaluate a 3D U-Net [41] for WM lesion segmentation, using patches of dimensions $96^{3}$. The split was performed at patient level to ensure that images from the same patient belong to the same split. We adopted a linear combination of normalized Dice [23] and blob [15] losses to tackle instance imbalance within a class and bias towards the occurrence of positive class [20]. Pre-processing steps included the registration of FLAIR images to MPRAGE space using the elastix toolbox [13, 27], N4 bias field inhomogeneity correction [34] and $z$-score intensity normalisation. The described model achieved a Dice score of 0.60 and a normalized Dice score of 0.71 on the test set, detecting 4868 WM lesion candidates out of the 3839 manually annotated.
2.2 Notations
@@
Following [8], we noted a discrete image as a $D$-dimensional function of the variable $\boldsymbol{v}=(v_{1},\dots,v_{D})\in\mathbb{Z}^{D}$, taking values $x[\boldsymbol{v}]\in\mathbb{R}$. A subset $\Gamma$ of $\mathbb{Z}^{D}$ was considered in practice for the spatial image domain with dimensions $N_{1}\times\cdots\times N_{D}$ as possible values for the index vector $\boldsymbol{v}\in\Gamma$. We also referred to the lesion domain $\Omega$ as a subset of the image domain with cardinality (i.e., number of voxels) $|\Omega|$, such that $\Omega\subset\Gamma\subset\mathbb{Z}^{D}$.
@@
Input images $x[\boldsymbol{v}]$ performing a forward pass through the network resulted in logits $y\left(x\right)[\boldsymbol{v}]\in\mathbb{R}$, where $\boldsymbol{v}\in\Gamma$ is a map of raw output values, which had the exact same dimensions as the input $x[\boldsymbol{v}]$ since we considered a segmentation task. We use the simplified notation $y[\boldsymbol{v}]$ when the input $x$ is unambiguous. This raw output was generally interpreted as a probability map (e.g., after Softmax), which was binarized using a threshold $t=0.3$ yielding the best normalized Dice score during validation. Then, each connected component (with a minimum volume size of $5mm^{3}$) of the binary map was considered as a specific WM lesion instance, forming the subdomain $\Omega$.
2.3 Gradient-based saliency maps
@@
During the backward pass and for each element of $y\left(x\right)[\boldsymbol{v}]$, the gradients with respect to all the elements of input $x[\boldsymbol{v}]$ can be computed. These gradients can be visualized as an image with same dimensions as the input and output, constituting a first method to construct saliency maps. This approach is usually referred to as vanilla gradients [28].
@@
2.3.1 Instance-level saliency (gradients)
@@
    Injecting a Gaussian noise $\mathcal{N}(0,\sigma)$ with standard deviation $\sigma$ to obtain $N$ noisy versions of the input,
@@
    Repeating steps 1-3 and combining these $N=50$ saliency maps ($\sigma=0.05$) to obtain a single one.
@@
Fig. 4: Overview of the proposed adaptation of SG to segmentation.
2.3.2 Quantitative saliency maps: maximum versus average aggregation
@@
2.4 Saliency maps based on Grad-CAM++
The second proposed method is based on Grad-CAM++ [5] that we first recall here. We generated, for a given layer of a network, a heatmap $M$ as a linear combination between weights $\{\omega^{k}\}^{K}_{k=1}$ and activation maps $\{A^{k}\}^{K}_{k=1}$. The activation maps of the selected layer may have different dimensions. In this case, the final heatmap is upsampled to the input dimensions. For the sake of simplicity, we defined the domain of the activation maps and that of the input image to be the same (i.e., activations from the last layer since we have a segmentation architecture): $\Gamma$, with values $A^{K}[\boldsymbol{v}]\in\mathbb{R}$. Following [36], to compute the gradients, we considered $y^{\prime}$ as the sum of logits $y[\boldsymbol{v}]$ higher than a threshold $t$, as in Eq.(5). Then, each weight $\omega^{k}$ is computed using the gradients of $y^{\prime}$ with respect to the $k^{th}$ activation map $A^{k}$ and a coefficient $\alpha^{k}[\boldsymbol{v}]$. We have
@@ 2.4 Saliency maps based on Grad-CAM++
where $\boldsymbol{v}$ and $\boldsymbol{v}^{\prime}$ correspond to a different indexing over the domain $\Gamma$. A graphical representation of this method is reported in Fig. 5.
Fig. 5: Overview of Grad-CAM++, generating a class-level explanation heatmap, similarly to [36], adapted for segmentation.
2.4.1 Instance-level saliency (Grad-CAM++)
@@
where $\alpha^{k}[\boldsymbol{v}]$ were obtained as in Eq. (7). An overview of the proposed adaptation of Grad-CAM++ to segmentation is illustrated in Fig. 6.
Fig. 6: Overview of the proposed adaptation of Grad-CAM++, generating an instance-level explanation heatmap.
@@ 2.5 Experiments
The described methodologies were applied to the entire test set, obtaining saliency maps for 3050 true positive (TP), 1818 false positive (FP), 789 false negative (FN), and 1010 true negative volumes (TN). TP and FP predictions were defined as having, respectively, a non-zero and zero overlap with ground truth (GT). FN predictions were considered as GT segmentations with zero overlap with the predicted lesion mask. To limit partial volume effects in these first three groups, a minimum volume size of $5mm^{3}$ was set for each connected component, using a connectivity of 2. TN examples were generated by randomly sampling ten spherical volumes ($93mm^{3}$) inside each patient’s brain and skull, excluding volumes intersecting GT and prediction masks. The size of the TN volumes was decided based on the average lesion volume in the GT masks of the test set.
@@
The information exploited by the U-Net during inference for the segmentation of specific WM lesions was assessed with several experiments. Throughout these tests, the max aggregation method (see Material and methods 2.3.2) was selected to be used, thanks to its ability to provide quantitative information. This only applies to the saliency based on SG.
@@
First (results in Section 3.2), the distribution of positive and negative values in saliency maps was observed to reveal potential patterns concerning recurring locations of positive/negative values with respect to $\Omega$. An analysis of these distributions allowed a statistical comparison between gradient values computed with respect to FLAIR and MPRAGE. As a consequence, it was possible to isolate the contribution of both input sequences to the prediction of single lesions. In this analysis, we discarded gradient values between -0.1 and 0.1 in order to focus on voxels with a higher attention level.
@@
Second (results in Section 3.3), we quantitatively compared the distribution of maximum and minimum saliency maps’ values, for all predictions categories (i.e., TP, FP, FN and TN) to investigate if absolute saliency values can be used to flag potential detection errors. Furthermore, a two-sided Mann Whitney U test [21] was run to statistically compare these groups.

### v3

@@ Abstract
In recent years, explainable methods for artificial intelligence (XAI) have tried to reveal and describe models’ decision mechanisms in the case of classification and even for segmentation. However, XAI methods for semantic segmentation and in particular for single specific instances (e.g. one given lesion among others of the same class in medical imaging) have yet to be developed to understand what drove the detection and contouring of the latter, which is crucial for all multi-lesional diseases.
We proposed instance-level explanation maps for semantic segmentation extending both SmoothGrad and Grad-CAM++ methods and yielding quantitative instance saliency for the former. The instance-level methods were applied to the segmentation of white matter lesions (WML), a magnetic resonance imaging (MRI) biomarker in multiple sclerosis (MS). 687 patients diagnosed with MS for a total of 4023 FLAIR and MPRAGE MRI scans were collected at the University Hospital of Basel, Switzerland. WM lesion masks were annotated by four expert clinicians on baseline and follow-up imaging. Three deep learning networks —a 3D U-Net, nnU-Net, and Swin UNETR— were trained and tested on these data (test normalized Dice score, respectively of 0.71, 0.78, 0.80; true positive rate of 79%, 78%, and 85%; false discovery rate of 37%, 38%, and 36%; false negative rate of 20%, 22%, and 14%), then saliency maps were computed.
Consistent with clinical practice, the proposed instance saliency maps revealed that the model relied more on FLAIR than MPRAGE to segment WMLs, with positive saliency values inside a lesion and negative in its neighborhood. FLAIR hyperintensity combined with healthy WM around the lesion border was required for their detection. Beyond the aforementioned sanity checks, we observed that peak values of the generated saliency maps presented distributions that differ significantly between TP, FN, FP and TN predictions, suggesting that the quantitative nature of the proposed saliency could be used to identify errors.
In conclusion, we introduced two XAI methods to generate quantitative instance-level explanations in semantic segmentation. The proposed XAI maps can be applied to any architecture and could serve as a basis to (i) improve model performance (e.g. reducing FPs), (ii) optimize their internal architecture (e.g. patch size), and (iii) justify the model’s decisions to the end users, which are contextualized to a specific lesion instance of interest.
@@
MRI, Multiple Sclerosis, XAI, deep learning, segmentation
Introduction
Multiple sclerosis (MS) is an autoimmune neurological disease, which affects people at a relatively young age, presenting a considerable impact on the quality of life [27]. One of the most important biomarkers in MS are white matter (WM) lesions reported on magnetic resonance imaging (MRI) [53]. The standard MRI sequences used for MS diagnosis and follow-up [47] are the fluid attenuated inversion recovery (FLAIR) and T1-weighted (T1-w) contrast, such as the magnetisation-prepared rapid gradient echo (MPRAGE) [20]. WM lesions generally appear hyperintense on FLAIR (except for cavitary lesions [4]), and a subset with greater demyelination and tissue damage appears hypointense on T1-weighted images [46].
@@ Introduction
    The “black box” nature of the models [5]. Since these methods contain many layers and millions of parameters it is hard to interpret and explain which are the drivers for a particular decision, i.e., which voxels were more important to identify and segment a given lesion of interest.
@@
    Insufficient clinical validation of the models, described in Spagnolo et al. [44].
To address the first issue, research in explainable AI (XAI) may play a decisive role. XAI has the potential to support trustworthy AI via better understanding (e.g., decision rules, biases) and optimization of DL models [25]. However, XAI for semantic segmentation and in particular for detecting and contouring single instances of interest has been little studied to date. Semantic segmentation is a computer vision task, where labels are associated with every pixel of an image. In the case of WM lesions, such as depicted in Fig.1a with four distinct lesion instances, the segmented plaques appear as disconnected volumes in the MRI. In general, objects may appear as connected, and can be segmented into separate entities through instance segmentation. Instance segmentation can extract supplementary information from the image, such as the number of objects of the same class (as lesion one to four in Fig.1b). Treating separate instances would be crucial to understand the mechanisms underlying automatic detection and segmentation of a given object of interest (e.g., a lesion in medical imaging or a person in a natural image) among other objects of the same class. This also applies to the case of MS lesions, where plaques can be subdivided based on their location in the brain, or the stage of the disease [23]. Specifically, considering separate instances would be important to generate instance-specific explanations not only for diagnosis at single time-points, but also for disease monitoring in follow-ups.
@@
Figure 1: (a) FLAIR MRI presenting WM lesions segmented as separate entities, and (b) example of instance segmentation inspired by Varatharasan et al.[50].
@@
An exhaustive review of XAI models and applications can be found in Saranya et al. [38]. For convolutional neural networks (CNN) in a classification scenario, a widely used ad-hoc method is pixel attribution (or saliency maps), in which pixels are colored based on their contribution to the classification. To this end, vanilla gradients [41] is a method based on forward and backward propagation through the network. First, an input image (for instance a 3D volume) is fed forward through the network and a class score is computed. Then, the gradient of the score with respect to the input — or a layer — is calculated backwards to form a map, which represents positive and negative contributions of input voxels to the classification of the image. Maps generated with this method are easy to compute and visualize, but also noisy and sensitive to small changes in the input [12]. To partially address these problems, Smilkov et al. [43] introduced a method called SmoothGrad (SG). A more stable output is obtained by feeding multiple noisy versions of the input image to the network and averaging the obtained saliency maps. Some recent applications can be found in Goh et al. [17] and Agarwal et al. [1].
@@
Another widely used XAI method for classification, is Grad-CAM [39]. The gradients of a class score — with respect to activation maps of a given layer — are spatially aggregated through global average pooling. This way, a weight is computed for each activation map, representing its relevance. These weights are then employed to calculate the visualization heatmap as a linear combination of the activation maps, followed by a ReLU to focus on features positively influencing the class score. However, this method presented a low accuracy when the image contained multiple instances of the same class. To overcome this, Chattopadhyay et al. [6] described Grad-CAM++, where the weights are obtained through a weighted average of the gradients.
SG and Grad-CAM++ can be seen as complementary methods. The first provides local-level information and works well in identifying the impact of multiple input channels (or modalities), since the gradients flow all the way back to the inputs. Yet, as mentioned above, it may be too sensitive to changes in the input. The second can generate more stable heatmaps, and probe specific layers of the network. However, it also bears its disadvantages: 1) the choice of the intermediate layer can be not trivial, 2) the potential presence of early skip-connections in the architecture would be disregarded, 3) stopping the backpropagation at an intermediate layer would make it impossible to determine the impact of the input on the output.
@@
Both algorithms were originally designed for classification tasks, such as in Rajpurkar et al. [37]. This means that, in general, the output $y$ of a multi-class classification network is a vector of $C$ scalar values, where $C$ is the number of classes. In the case of semantic segmentation, $y$ is a set of $C$ tensors (2D or 3D images) containing segmentation scores for each class (Fig. 2).
@@
Figure 2: Input-output dimensions for a classification (top) and a semantic segmentation (bottom) task.
@@
A straightforward way to adapt explainable methods to semantic segmentation would be, for a given class $c$, to aggregate all the spatial predictions $y^{c}$ into a single scalar (e.g., a summation) and compute its gradients. However, this approach yields confusing and hardly interpretable maps, where the relevance of input voxels to the segmentation scores of all output voxels (even those not segmented as part of class $c$) are merged together, as reported in Fig. 3. This makes it more complicated to understand and determine which part of the output segmentation is influenced by which region of the input. That is especially important when there are multiple objects belonging to the same class, so that input intensity of neighboring voxels or objects may impact the segmentation. Indeed, these results do not provide any kind of explanation to the segmentation of a particular instance of the considered class $c$.
Figure 3: (a) The output of a semantic segmentation network showing several instances of the considered class. SmoothGrad (b) and Grad-CAM (c) applied to all the spatial predictions. (d) How can we explain the segmentation of a particular lesion of interest (e.g., the yellow instance)?
Recent works [31, 3] clearly state that saliency maps were not initially developed for segmentation tasks and, therefore, no explicit methods are currently available to do so. In Arun et al. [3], this statement was used to warn the readers about potential problems related to the misuse of XAI in clinical practice, and proposed pixel-level metrics to evaluate saliency maps. Singh et al. [42] applied saliency maps to classification networks, while adopting uncertainty maps to assess the confidence of their skin lesion segmentation method.
First steps towards the use of XAI in segmentation were taken in the work of Wickstrom et al. [52]. They used guided backpropagation [41] to generate saliency maps for the explanation of colorectal polyp segmentation. Their saliency maps were obtained by aggregating all the positive spatial predictions. Similarly, Vinogradova et al. [51] generated heatmaps by using the first version of Grad-CAM [39], and aggregating only output voxels segmented as part of a target class $c$. The authors employed a U-Net and the dataset Cityscapes [10] to perform semantic segmentation.
@@
The quantitative characteristic of the saliency method described in Section Quantitative saliency maps: maximum versus average aggregation was exploited in Spagnolo et al. [45], using the same MRI data. Radiomic features extracted from the proposed instance-level XAI maps were fed to a simple logistic regression model to refine the classification of TP and FP examples. In that study, a bootstrapping approach with 1000 iterations was used to compute the 95% confidence intervals, highlighting an F1 score relative improvement on the test set.
Methods
Dataset and models
687 patients diagnosed with MS for a total of 4023 FLAIR and MPRAGE MRI scans (age=45.2±12.2, 433 females, SIEMENS Avanto/Espree/Symphony 1.5T and Prisma/Skyra/Verio/MAGNETOM Vida 3T, $1mm$ isotropic, Expanded Disability Status Scale median of 2.5 [0-9]) were collected at the University Hospital of Basel, Switzerland [15]. The study received ethical approval by the local independent ethics committee (Ethikkommission Nordwest- und Zentralschweiz, EKNZ), all patients provided written informed consent, and all methods were carried out by relevant guidelines and regulations.
The image size of both MR contrasts is (192 × 240 × 256), which corresponds to a volume of (192 × 240 × 256) $mm^{3}$. Table LABEL:supptable1 in the Supplementary Material reports data information for each MR system, such as age, sex and number of visits. For consistency reasons, patients were mainly scanned first using Avanto 1.5T and then Skyra 3T, operating other MR systems only in case of unavailability of the two mentioned.
@@ Methods
WM lesion masks from baseline and follow-ups were semi-automatically annotated by four expert clinicians (>5 years of experience), independently and without consensus. The binary lesion masks corrected by the experts were generated by a U-Net variant described in La Rosa et al. (2020)[28] (since that variant was trained on FLAIR and MP2RAGE data, we retrained it on separate proprietary FLAIR and MPRAGE data before the inference). Data were randomly split into training, validation and test sets (containing 560, 90 and 37 patients with 3369, 553 and 101 scans, respectively; training, validation and hold-out test set’s mean lesions number of $52.0\pm 36.3$, $56.2\pm 35.3$, and $42.3\pm 21.4$ per patient) to train and evaluate a 3D U-Net [7] variant (different from the one in La Rosa et al [28]), an nnU-Net [21], and a Swin UNETR [19] for WM lesion segmentation, using patches of dimensions $96^{3}$ to ensure the inclusion of at least part of the brain in every patch. The split was performed at patient level to ensure that images from the same patient belong to the same split. We adopted a linear combination of normalized Dice [36] and blob [26] losses to tackle instance imbalance within a class and bias towards the occurrence of positive class [32]. Pre-processing steps included the registration of FLAIR images to MPRAGE space using the elastix toolbox [24, 40], N4 bias field inhomogeneity correction [49] and $z$-score intensity normalisation.
Notations
@@
Following Depeursinge et al. (2020) [13], we noted a discrete image as a $D$-dimensional function of the variable $\boldsymbol{v}=(v_{1},\dots,v_{D})\in\mathbb{Z}^{D}$, taking values $x[\boldsymbol{v}]\in\mathbb{R}$. A subset $\Gamma$ of $\mathbb{Z}^{D}$ was considered in practice for the spatial image domain with dimensions $N_{1}\times\cdots\times N_{D}$ as possible values for the index vector $\boldsymbol{v}\in\Gamma$. We also referred to the lesion domain $\Omega$ as a subset of the image domain with cardinality (i.e. number of voxels) $|\Omega|$, such that $\Omega\subset\Gamma\subset\mathbb{Z}^{D}$.
@@
Input images $x[\boldsymbol{v}]$ performing a forward pass through the network resulted in logits $y\left(x\right)[\boldsymbol{v}]\in\mathbb{R}$, where $\boldsymbol{v}\in\Gamma$ is a map of raw output values, which had the exact same dimensions as the input values $x[\boldsymbol{v}]$ since we considered a segmentation task. We use the simplified notation $y[\boldsymbol{v}]$ when the input $x$ is unambiguous. This raw output was generally interpreted as a probability map (e.g., after Softmax), which was binarized using a threshold $t=0.3$ yielding the best normalized Dice score during validation. Various thresholds $t$ were explored, ranging from 0.1 to 0.5, and the one selected scored about 2% better than the second best. Then, each connected component of the binary map was considered as a specific WM lesion instance, forming $\Omega$.
Gradient-based saliency maps
@@
During the backward pass and for each voxel of $y\left(x\right)[\boldsymbol{v}]$, the gradients with respect to all the voxels of input values $x[\boldsymbol{v}]$ can be computed. These gradients can be visualized as an image with same dimensions as the input and output, constituting a first method to construct saliency maps. This approach is usually referred to as vanilla gradients [41].
@@
Instance-level saliency (gradients)
@@
    Injecting a Gaussian noise $\mathcal{N}(0,\sigma)$ with standard deviation $\sigma$ to obtain a noisy version of the input,
@@
    Repeating steps 1-3 and combining $N=50$ saliency maps ($\sigma=0.05$) from $N$ noisy versions to obtain a single one.
@@
Figure 4: Overview of the proposed adaptation of SG to segmentation.
Quantitative saliency maps: maximum versus average aggregation
@@
Saliency maps based on Grad-CAM++
The second proposed method is based on Grad-CAM++ [6]. We generated, for a given layer of a network, a heatmap $M$ as a linear combination between weights $\{\omega^{k}\}^{K}_{k=1}$ and activation maps $\{A^{k}\}^{K}_{k=1}$. The activation maps of the selected layer may have different dimensions. In this case, the final heatmap is upsampled to the input dimensions. For the sake of simplicity, we defined the domain of the activation maps and that of the input image to be the same (i.e. activations from the last layer since we have a segmentation architecture): $\Gamma$, with values $A^{K}[\boldsymbol{v}]\in\mathbb{R}$. Following Vinogradova et al. [51], to compute the gradients, we considered $y^{\prime}$ as the sum of logits $y[\boldsymbol{v}]$ higher than a threshold $t$, as in Eq.(5). Then, each weight $\omega^{k}$ is computed using the gradients of $y^{\prime}$ - with respect to the $k^{th}$ activation map $A^{k}$ - and a coefficient $\alpha^{k}[\boldsymbol{v}]$. We have
@@
where $\boldsymbol{v}$ and $\boldsymbol{v}^{\prime}$ correspond to a different indexing over the domain $\Gamma$. The above eqs.˜4, 5, 6 and 7 are derived from Chattopadhyay et al. (2018) [6]. A graphical representation of this method is reported in Fig. 5.
Figure 5: Overview of Grad-CAM++, generating a class-level explanation heatmap, similarly to Vinogradova et al. [51], adapted for segmentation.
Instance-level saliency (Grad-CAM++)
@@
where $\alpha^{k}[\boldsymbol{v}]$ were obtained as in Eq. (7). An overview of the proposed adaptation of Grad-CAM++ to segmentation is illustrated in Fig. 6. Comparing Figs 5 and 6, it can be observed that the output map of the first is a valuable explanation for the entire lesion class, combining the attribution of voxels from different lesions in the same map. Conversely, the output of the second separates the attribution of voxels from single lesions to obtain a more meaningful map, in cases where each instance should be treated independently.
Figure 6: Overview of the proposed adaptation of Grad-CAM++, generating an instance-level explanation heatmap.
@@
The described methodologies were applied to all the three selected architectures (U-Net, nnU-Net and Swin UNETR), obtaining a collection of saliency maps. The XAI maps presented below were computed using the U-Net for a total of 3050 true positive (TP), 1818 false positive (FP), 789 false negative (FN), and 1010 true negative volumes (TN). Examples of saliency maps derived from nnU-Net and Swin UNETR are included in the Supplementary Material. TP and FP predictions were defined as having, respectively, a non-zero and zero overlap with ground truth (GT). FN predictions were considered as GT segmentations with zero overlap with the predicted lesion mask. TN examples were generated by randomly sampling ten spherical volumes ($93mm^{3}$) inside each patient’s brain and skull, excluding volumes intersecting GT and prediction masks. The size of the TN volumes was decided based on the average lesion volume in the GT masks of the test set. In this study, when mentioning TP, FP, FN and TN we refer to examples, i.e. volumes, not voxels. While TN voxels represent most of the image, we tackle class imbalance by undersampling, randomly selecting ten spherical volumes per patient.
As for the minimum considered lesion size, the McDonald’s criteria recommend a minimum in-plane diameter of 3 $mm$ based on old 2D sequences [35, 18], and in literature there is no clear consensus. As a trade off between limiting partial volume effects and the number of false positives [16], a minimum volume size of $5mm^{3}$ was set for each connected component, using a connectivity of 18. Several publications have selected a similar minimum lesion volume, for example: 1) De Rosa et al. (2024)[11] considered 0.005 $ml$ (5 $mm^{3}$); 2) Fartaria et al. (2019)[16] used 0.006 $ml$; 3) Jain et al. (2016)[22] removed candidates with a volume smaller than 0.005 $ml$.
@@
The information exploited by the U-Net during inference for the segmentation of specific WM lesions was assessed with several experiments. Throughout these tests, the max aggregation method (see Section Quantitative saliency maps: maximum versus average aggregation) was selected to be used, due to its ability to provide quantitative information. This only applies to the saliency based on SG.
@@
First (results in Section Assessing the contributions of input MR sequences using gradient-based saliency maps), the distribution of positive and negative values in saliency maps was observed to reveal potential patterns concerning recurring locations of positive/negative values with respect to $\Omega$. An analysis of these distributions allowed a statistical comparison between gradient values computed with respect to FLAIR and MPRAGE. As a consequence, it was possible to isolate the contribution of both input sequences to the prediction of single lesions. In this analysis, we discarded gradient values between -0.1 and 0.1 in order to focus on voxels with a higher attention level.
@@
Second (results in Section Statistical distribution of gradient values for TPs, FPs, FNs and TNs), we quantitatively compared the distribution of maximum and minimum saliency maps’ values, for all predictions categories (i.e. TP, FP, FN and TN) to investigate if absolute saliency values can be used to flag potential detection errors. Furthermore, a two-sided Mann Whitney U test [33] was run to statistically compare these groups.
Third (results in Section Sanity checks), we tested the saliency method’s behavior in two cases: (a) positioning a domain $\Omega$ in a region presenting no MS lesions (i.e. healthy WM); (b) considering a domain $\Omega$ of a single voxel, located at the center of mass of a true lesion. For both cases, we show the generated instance-level saliency and its range of values.